# On multi-graded Proj schemes

### v3

@@ 3.2 Open subschemes defined by quasi-relevant elements
Recall the notion of quasi-$M$-relevant element of $A$ from Definition 2.11. If $a\in A$ is $M$-quasi-relevant, then we can write $a=\sum_{i=1}^{n}f_{i}$ where each $f_{i}$ is $M$-relevant of degree $\deg(a)$, and by Proposition 2.10(4) we have an open immersion
@@
Assume that $a=\sum_{i=1}^{n}f_{i}=\sum_{j=1}^{m}g_{j}$ where the $f_{i}$’s and the $g_{j}$’s are relevant, all of the same degree. What we have to prove is that the compositions
@@
By (3.3), ${\operatorname{Proj}}^{M}(A)$ is covered by the open subschemes $D_{\dagger}(f)$ where $f$ is relevant. By assumption, for such $f$ there exists $k\in\mathbb{Z}_{>0}$ such that $f^{k}=\sum_{j=1}^{r}b_{j}\cdot a_{i_{j}}$ for some elements $b_{j}\in A$ and some indices $i_{j}\in I$. Since $f$ is homogeneous, here we can assume that for any $j$ the element $b_{j}$ is homogeneous, with $\deg(b_{j})+\deg(a_{i_{j}})=\deg(f)$. Then we have
@@
3.3 Some basic properties
@@
as open subschemes of ${\operatorname{Proj}}^{M}_{\Psi({\mathcal{F}})}(B)$. On the other hand, by definition ${\operatorname{Proj}}^{M}_{\Psi({\mathcal{F}})}(B)$ is covered by the open subschemes $D_{\dagger}(\Psi(T))$ when $T$ runs over $\mathcal{F}$, so that ${\operatorname{Proj}}^{M}_{\Psi({\mathcal{F}})}(B)\times_{{\operatorname{Proj}}_{{\mathcal{F}}}^{M}(A)}D_{\dagger}(S)$ is covered by the open subschemes
$D_{\dagger}(\Psi(T))\times_{{\operatorname{Proj}}_{{\mathcal{F}}}^{M}(A)}D_{\dagger}(S).$
Now the morphism $D_{\dagger}(\Psi(T))\to{\operatorname{Proj}}_{{\mathcal{F}}}^{M}(A)$ factors through $D_{\dagger}(T)$, so that
@@
Using [StP, Tag 01S8], the equalities (3.6) imply that our morphism is affine. ∎
@@
In the setting of Proposition 3.12, assume that $\Psi$ is surjective. Then we have $\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)=\mathrm{Proj}^{M}(B)$, and the morphism
${\operatorname{Proj}}^{M}(B)\to{\operatorname{Proj}}^{M}(A)$
is a closed immersion.
@@
Let $S\in{\mathcal{F}}_{B}$. By definition, there exist $s_{1},\dots,s_{r}\in\underline{S}$ whose degrees generate a subgroup of $M$ of finite index. Let $S^{\prime}\subset B$ be the homogenenous multiplicative subset generated by $s_{1},\dots,s_{r}$, and let $S^{\prime\prime}=S\cdot S^{\prime}$. Then $S^{\prime}$ and $S^{\prime\prime}$ are relevant homogeneous multiplicative subsets of $B$, and by Remark 3.2(3) we have $D_{\dagger}(S)=D_{\dagger}(S^{\prime\prime})$. We therefore have an embedding $D_{\dagger}(S)\to D_{\dagger}(S^{\prime})$. Now, choosing for each $i\in\{1,\dots,r\}$ a homogeneous preimage $t_{i}$ of $s_{i}$ in $A$ and denoting by $\tilde{S}\subset A$ the homogeneous multiplicative subset generated by $t_{1},\dots,t_{r}$, it is clear that $\tilde{S}$ is relevant and that $\Psi(\tilde{S})=S^{\prime}$. Hence $D_{\dagger}(S^{\prime})\subset\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)$, which implies that $D_{\dagger}(S)$ is also contained in $\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)$. Since $S$ was arbitrary, this proves the equality $\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)=\mathrm{Proj}^{M}(B)$.
@@
is a closed immersion. This follows from (3.6), since the morphism $A_{(f)}\to B_{(\Psi(f))}$ is surjective. ∎
@@
The identification is provided by the following equalities:
@@
Here the first equality follows from (3.3). The second equality follows from a basic property of fiber products of schemes, see [StP, Tag 01JS]. The third equality follows from the obvious isomorphism of $(M\times M^{\prime})$-graded rings
$A_{f}\otimes_{R}A^{\prime}_{f^{\prime}}\cong(A\otimes_{R}A^{\prime})_{f\otimes f^{\prime}}$
(for $f\in A$ and $f^{\prime}\in A^{\prime}$ homogeneous) by restriction to the components of degree $(0,0)$.
To conclude we have to explain the fourth equality. Let ${\mathcal{F}}_{A\otimes_{R}A^{\prime}}^{\otimes}\subset{\mathcal{F}}_{A\otimes_{R}A^{\prime}}$ be the subset consisting of homogeneous multiplicative subsets of the form $(f\otimes f^{\prime})^{\mathbb{N}}$ with $f\in A$ $M$-relevant and $f^{\prime}\in A^{\prime}$ $M^{\prime}$-relevant. (Here we use once again the notation introduced in Definition 2.3(3).) The desired equality will follow from the equality
${\operatorname{Proj}}_{{\mathcal{F}}_{\otimes}}^{M\times M^{\prime}}(A\otimes_{R}A^{\prime})={\operatorname{Proj}}^{M\times M^{\prime}}(A\otimes_{R}A^{\prime}),$
which can be justified as follows. By definition, the left-hand side is an open subset of the right-hand side. To prove that this open subset is all of ${\operatorname{Proj}}^{M\times M^{\prime}}(A\otimes_{R}A^{\prime})$, by (3.3) it suffices to prove that it contains the open subscheme attached to any relevant element $x$ in $A\otimes_{R}A^{\prime}$. Now if $x\in A\otimes_{R}A^{\prime}$ is relevant, by Proposition 2.5(1) there exist a positive integer $k$, a finite set $J$ and for any $j\in J$ an $M$-relevant element $s_{j}$ in $A$ and an $M^{\prime}$-relevant element $s^{\prime}_{j}$ in $A$ such that $x^{k}=\sum_{j\in J}s_{j}\otimes s^{\prime}_{j}$. Then we have
$D_{\dagger}(x)=D_{\dagger}(x^{k})\subset\bigcup_{j\in J}D_{\dagger}(s_{j}\otimes s^{\prime}_{j}),$
which finishes the proof. ∎
@@
${\operatorname{Proj}}(A\otimes_{R}R^{\prime})\cong\mathrm{Proj}(A)\times_{\mathrm{Spec}(R)}\mathrm{Spec}(R^{\prime}).$
@@
In view of Example 3.8, the statement is the special case of Proposition 3.15 corresponding to $M^{\prime}=\{0\}$. ∎
@@
Let $S\in{\mathcal{F}}_{A}$, and let $S_{R^{\prime}}$ be its image in $A\otimes_{R}R^{\prime}$. Then it is clear that under the isomorphism of Corollary 3.17 the open subscheme $D_{\dagger}(S_{R^{\prime}})\subset{\operatorname{Proj}}(A\otimes_{R}R^{\prime})$ identifies with $D_{\dagger}(S)\times_{\mathrm{Spec}(R)}\mathrm{Spec}(R^{\prime})$.
@@
Let $R$ be a commutative ring and $N$ be a commutative cancellative monoid such that the group $M=N-N$ is finitely generated. Let $R[N]:=\bigoplus_{m\in N}R\cdot X^{m}$ be the $R$-algebra of the monoid $N$, with its natural $M$-grading. The homogeneous elements in this ring are those of the form $rX^{m}$ with $r\in R$ and $m\in N$; such an element is relevant if and only if $N\cap(m-N)$ generates $M$. For each such element we have $(R[N])_{(rX^{m})}=R_{r}$; therefore we have $\mathrm{Proj}(R[N])=\mathrm{Spec}(R)$.
@@
We use the description of these schemes using (3.3). If $f$ is an $M$-relevant element in $A$, then by (3.2) we have $A_{(f)}=A_{(f^{d})}$ for all $d\in{\mathbb{Z}}_{>0}$. If $n$ is the index of $M^{\prime}$ in $M$, then $f^{n}$ belongs to $A^{\prime}$ and is $M^{\prime}$-relevant in $A^{\prime}$; moreover we have $A_{(f)}=A_{(f^{n})}=A^{\prime}_{(f^{n})}$. This shows that ${\operatorname{Proj}}^{M}(A)$ identifies with an open subscheme of ${\operatorname{Proj}}^{M^{\prime}}(A^{\prime})$. Reciprocally, if $f\in A^{\prime}$ is $M^{\prime}$-relevant, then $f$ is an $M$-relevant element of $A$; it follows that this open subscheme is the whole of ${\operatorname{Proj}}^{M^{\prime}}(A^{\prime})$.
@@
Remark 3.21.
Note that Proposition 3.20 is not true if $M/M^{\prime}$ is infinite. For example, let $R$ be a ring, which we endow with the ${\mathbb{Z}}$-grading where $R$ is in degree 00. One can also consider $R$ as a $\{0\}$-graded. There is no ${\mathbb{Z}}$-relevant element in $R$ so ${\operatorname{Proj}}^{\mathbb{Z}}(R)=\varnothing$; however by Example 3.8 we have ${\operatorname{Proj}}^{\{0\}}(R)={\operatorname{Spec}}(R)$.
@@
    Let $A$ and $B$ be commutative $M$-graded rings, and assume that there exists a subgroup $M^{\prime}\subset M$ such that $M/M^{\prime}$ is finite and the subrings of $A$ and $B$ consisting of their homogeneous components associated with elements in $M^{\prime}$ are isomorphic (as $M^{\prime}$-graded rings). Then we have
    ${\operatorname{Proj}}^{M}(A)\cong{\operatorname{Proj}}^{M}(B).$
@@ 3.4 Relative Proj schemes
Proposition 3.17 allows us to extend the construction of Proj schemes to a relative setting, as follows. Let $X$ be a scheme, $M$ be a finitely generated abelian group, and $\mathcal{A}$ be an $M$-graded quasi-coherent $\mathcal{O}_{X}$-algebra.
Construction 3.23 ((Relative Proj)).
@@
3.5 Proj, blowups and dilatations
Let $A$ be a ring and $\{J_{i}:i\in I\}$ be ideals of $A$. Recall the ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A=\bigoplus_{\nu\in\mathbb{N}_{I}}J_{i}^{\nu}$ (cf. Example 2.13 and [Ma23b, Definition 2.28]), which is $\mathbb{Z}_{I}$-graded. We assume now that $I$ if finite. Any family $(a_{i}:i\in I)$ of elements of $A$ such that $a_{i}\in J_{i}$ for each $i\in I$ can be considered as a relevant family in the $\mathbb{Z}_{I}$-graded ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$, so we have an associated affine open subscheme
$D_{\dagger}(a_{i}:i\in I)\subset{\operatorname{Proj}}^{\mathbb{Z}_{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$
and this open subscheme identifies with the dilatation ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$ (cf. Example 2.13 and [Ma23b, Fact 2.29]).
The goal of the present subsection is to study related connections and compatibilities between dilatations and these specific Proj schemes, and to show that this allows to easily provide a proof of the existence of (projective) blowups, possibly with multiple centers. We start with a multi-centered version of [StP, Tag 0804].

### v4

@@ 3.3 Open subschemes defined by quasi-relevant elements
Recall the notion of quasi-$M$-relevant element of $A$ from Definition 2.11. If $a\in A$ is a nonzero $M$-quasi-relevant element, then we can write $a=\sum_{i=1}^{n}f_{i}$ where each $f_{i}$ is a nonzero $M$-relevant element of degree $\deg(a)$, and by Proposition 2.10(4) we have an open immersion
@@
Assume that $a=\sum_{i=1}^{n}f_{i}=\sum_{j=1}^{m}g_{j}$ where the $f_{i}$’s and the $g_{j}$’s are nonzero and relevant, all of the same degree. What we have to prove is that the compositions
@@
Of course we can assume that the $a_{i}$’s are nonzero. By (3.3), ${\operatorname{Proj}}^{M}(A)$ is covered by the open subschemes $D_{\dagger}(f)$ where $f$ is relevant and not nilpotent. By assumption, for such $f$ there exists $k\in\mathbb{Z}_{>0}$ such that $f^{k}=\sum_{j=1}^{r}b_{j}\cdot a_{i_{j}}$ for some nonzero elements $b_{j}\in A$ and some indices $i_{j}\in I$. Since $f$ is homogeneous, here we can assume that for any $j$ the element $b_{j}$ is homogeneous, with $\deg(b_{j})+\deg(a_{i_{j}})=\deg(f)$. Then we have
@@
  1. 1.
    Alternatively, the corollary can be proved by remarking that our assumptions imply that $V(A_{\dagger})=\bigcap_{i\in I}V(A\cdot a_{i})$, so that ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})=\bigcup_{i\in I}D(a_{i})$, and then using the map (3.4).
  2. 2.
@@
3.4 More basic properties
@@
as open subschemes of ${\operatorname{Proj}}^{M}_{\Psi({\mathcal{F}})}(B)$. On the other hand, by definition ${\operatorname{Proj}}^{M}_{\Psi({\mathcal{F}})}(B)$ is covered by the open subschemes $D_{\dagger}(\Psi(T))$ where $T$ runs over $\mathcal{F}$, so that ${\operatorname{Proj}}^{M}_{\Psi({\mathcal{F}})}(B)\times_{{\operatorname{Proj}}_{{\mathcal{F}}}^{M}(A)}D_{\dagger}(S)$ is covered by the open subschemes $D_{\dagger}(\Psi(T))\times_{{\operatorname{Proj}}_{{\mathcal{F}}}^{M}(A)}D_{\dagger}(S)$. Now the morphism $D_{\dagger}(\Psi(T))\to{\operatorname{Proj}}_{{\mathcal{F}}}^{M}(A)$ factors through $D_{\dagger}(T)$, so that
@@
The equalities (3.6) imply that our morphism is affine, see e.g. [StP, Tag 01S8]. ∎
@@
In the setting of Proposition 3.12, assume that $\Psi:A\to B$ is surjective. Then we have $\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)=\mathrm{Proj}^{M}(B)$, and the canonical morphism ${\operatorname{Proj}}^{M}(B)\to{\operatorname{Proj}}^{M}(A)$ is a closed immersion.
@@
Let $S\in{\mathcal{F}}_{B}$. By definition, there exist $s_{1},\dots,s_{r}\in\underline{S}$ whose degrees generate a subgroup of $M$ of finite index. Let $S^{\prime}\subset B$ be the homogenenous multiplicative subset generated by $s_{1},\dots,s_{r}$, and let $S^{\prime\prime}=S\cdot S^{\prime}$. Then $S^{\prime}$ and $S^{\prime\prime}$ are finitely generated relevant homogeneous multiplicative subsets of $B$, and by Remark 3.2(3) we have $D_{\dagger}(S)=D_{\dagger}(S^{\prime\prime})$. We therefore have an embedding $D_{\dagger}(S)\to D_{\dagger}(S^{\prime})$. Now, choosing for each $i\in\{1,\dots,r\}$ a homogeneous preimage $t_{i}$ of $s_{i}$ in $A$ and denoting by $\tilde{S}\subset A$ the homogeneous multiplicative subset generated by $t_{1},\dots,t_{r}$, it is clear that $\tilde{S}$ is relevant and that $\Psi(\tilde{S})=S^{\prime}$. Hence $D_{\dagger}(S^{\prime})\subset\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)$, which implies that $D_{\dagger}(S)$ is also contained in $\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)$. Since $S$ was arbitrary, this proves the equality $\mathrm{Proj}^{M}_{\Psi(\mathcal{F}_{A})}(B)=\mathrm{Proj}^{M}(B)$.
@@
is a closed immersion. This follows from (3.6), since the morphism $A_{(f)}\to B_{(\Psi(f))}$ is clearly surjective. ∎
@@
The identification is provided by the following equalities (to be explained below):
@@
Here the first equality follows from (3.3). The second equality follows from a basic property of fiber products of schemes, see [StP, Tag 01JS]. The third equality follows from the obvious isomorphism of $(M\times M^{\prime})$-graded rings $A_{f}\otimes_{R}A^{\prime}_{f^{\prime}}\cong(A\otimes_{R}A^{\prime})_{f\otimes f^{\prime}}$ (for $f\in A$ and $f^{\prime}\in A^{\prime}$ homogeneous) by restriction to the components of degree $(0,0)$.
To conclude we have to explain the fourth equality. Let ${\mathcal{F}}_{A\otimes_{R}A^{\prime}}^{\otimes}\subset{\mathcal{F}}_{A\otimes_{R}A^{\prime}}$ be the subset consisting of homogeneous multiplicative subsets of the form $(f\otimes f^{\prime})^{\mathbb{N}}$ with $f\in A$ $M$-relevant and $f^{\prime}\in A^{\prime}$ $M^{\prime}$-relevant. The desired equality will follow from the equality
${\operatorname{Proj}}_{{\mathcal{F}}_{A\otimes_{R}A^{\prime}}^{\otimes}}^{M\times M^{\prime}}(A\otimes_{R}A^{\prime})={\operatorname{Proj}}^{M\times M^{\prime}}(A\otimes_{R}A^{\prime}),$
which can be justified as follows. By definition, the left-hand side is an open subset of the right-hand side. To prove that this open subset is all of ${\operatorname{Proj}}^{M\times M^{\prime}}(A\otimes_{R}A^{\prime})$, by (3.3) it suffices to prove that it contains the open subscheme attached to any relevant element $x$ in $A\otimes_{R}A^{\prime}$. Now if $x\in A\otimes_{R}A^{\prime}$ is relevant, by Proposition 2.5(1) there exist a positive integer $k$, a finite set $J$ and for any $j\in J$ an $M$-relevant element $s_{j}$ in $A$ and an $M^{\prime}$-relevant element $s^{\prime}_{j}$ in $A^{\prime}$ such that $x^{k}=\sum_{j\in J}s_{j}\otimes s^{\prime}_{j}$. Then we have
$D_{\dagger}(x)=D_{\dagger}(x^{k})\subset\bigcup_{j\in J}D_{\dagger}(s_{j}\otimes s^{\prime}_{j})$
by (3.2) and the constructions of §3.3, which finishes the proof. ∎
@@
${\operatorname{Proj}}^{M}(A\otimes_{R}R^{\prime})\cong{\operatorname{Proj}}^{M}(A)\times_{\mathrm{Spec}(R)}\mathrm{Spec}(R^{\prime}).$
@@
By Example 3.3, the statement is the special case of Proposition 3.15 when $M^{\prime}=\{0\}$. ∎
@@
Let $S\in{\mathcal{F}}_{A}$, and let $S_{R^{\prime}}$ be its image in $A\otimes_{R}R^{\prime}$. Then it is clear that under the isomorphism of Corollary 3.17 the open subscheme $D_{\dagger}(S_{R^{\prime}})\subset{\operatorname{Proj}}^{M}(A\otimes_{R}R^{\prime})$ identifies with $D_{\dagger}(S)\times_{\mathrm{Spec}(R)}\mathrm{Spec}(R^{\prime})$.
@@
Let $R$ be a commutative ring and $N$ be a commutative cancellative monoid such that the group $M=N^{\mathrm{gp}}$ is finitely generated. (See §1.6.3 for the terminology and notation.) Let $R[N]:=\bigoplus_{m\in N}R\cdot X^{m}$ be the $R$-algebra of the monoid $N$, with its natural $M$-grading. The homogeneous elements in this ring are those of the form $rX^{m}$ with $r\in R$ and $m\in N$; such an element is relevant if and only if $N\cap(m-N)$ generates $M$. For each such element we have $(R[N])_{(rX^{m})}=R_{r}$; therefore we have $\mathrm{Proj}(R[N])=\mathrm{Spec}(R)$.
@@
We use the description of these schemes using (3.3). If $f$ is an $M$-relevant element in $A$, then as in (3.2) we have $A_{(f)}=A_{(f^{d})}$ for all $d\in{\mathbb{Z}}_{>0}$. If $n$ is the index of $M^{\prime}$ in $M$, then $f^{n}$ belongs to $A^{\prime}$ and is $M^{\prime}$-relevant in $A^{\prime}$; moreover we have $A_{(f)}=A_{(f^{n})}=A^{\prime}_{(f^{n})}$. This shows that ${\operatorname{Proj}}^{M}(A)$ identifies with an open subscheme of ${\operatorname{Proj}}^{M^{\prime}}(A^{\prime})$. Reciprocally, if $f\in A^{\prime}$ is $M^{\prime}$-relevant, then $f$ is an $M$-relevant element of $A$; it follows that this open subscheme is the whole of ${\operatorname{Proj}}^{M^{\prime}}(A^{\prime})$.
@@
    Let $A$ and $B$ be commutative $M$-graded rings, and assume that there exists a subgroup $M^{\prime}\subset M$ such that $M/M^{\prime}$ is finite and the subrings of $A$ and $B$ consisting of their homogeneous components associated with elements in $M^{\prime}$ are isomorphic (as $M^{\prime}$-graded rings). Then we have ${\operatorname{Proj}}^{M}(A)\cong{\operatorname{Proj}}^{M}(B)$.
@@ 3.5 Relative Proj schemes
Proposition 3.17 allows us to extend the construction of Proj schemes to a relative setting, as follows. Let $X$ be a scheme, $M$ be a finitely generated abelian group, and $\mathcal{A}$ be an $M$-graded quasi-coherent $\mathcal{O}_{X}$-algebra. Note that for any quasi-compact open subscheme $U\subset X$ the algebra $\Gamma(U,\mathcal{A})$ admits a canonical $M$-grading by [StP, Tag 01AI].
Construction 3.22 ((Relative Proj)).
@@
3.6 Proj schemes, blowups and dilatations
Let $A$ be a ring and $\{J_{i}:i\in I\}$ be a family of ideals of $A$ parametrized by a finite set $I$. Recall the ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A=\bigoplus_{\nu\in\mathbb{N}^{I}}J_{i}^{\nu}$ (see Example 2.13 and [Ma23b, Definition 2.28]), which is $\mathbb{Z}^{I}$-graded. Any family $\{a_{i}:i\in I\}$ of elements of $A$ such that $a_{i}\in J_{i}$ for each $i\in I$ can be considered as a relevant family in the $\mathbb{Z}^{I}$-graded ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$, so that we have an associated affine open subscheme
$D_{\dagger}(\{a_{i}:i\in I\})\subset{\operatorname{Proj}}^{\mathbb{Z}^{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A),$
and this open subscheme identifies with the dilatation ring ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$ by Example 2.13.
The goal of the present subsection is to study related connections and compatibilities between dilatations and these specific Proj schemes, and to show that this allows to easily provide a proof of the existence of (projective) blowups, possibly with multiple centers. This section expands [Ma23b, § 3.8]. We start with a multi-centered version of [StP, Tag 0804].